OpenStudy (chely2013):

how do i find the derivative of the natural log of x squared? [ln(x^2)]

OpenStudy (anonymous):

d/dx(lnx2) = d/dx2(lnx2) . dx2/dx = 1/x2 . 2x = 2/x

OpenStudy (anonymous):

you got to use chain rule

OpenStudy (anonymous):

or simply power/x i.e 2/x

OpenStudy (chely2013):

thank you so much!

OpenStudy (anonymous):

welcome very much